### v1

# MagicColor: Multi-Instance Sketch Colorization

Abstract
We present MagicColor, a diffusion-based framework for multi-instance sketch colorization. The production of multi-instance 2D line art colorization adheres to an industry-standard workflow, which consists of three crucial stages: the design of line art characters, the coloring of individual objects, and the refinement process. The artists are required to repeat the process of coloring each instance one by one, which is inaccurate and inefficient. Meanwhile, current generative methods fail to solve this task due to the challenge of multi-instance pair data collection. To tackle these challenges, we incorporate three technical designs to ensure precise character detail transcription and achieve multi-instance sketch colorization in a single forward. Specifically, we first propose the self-play training strategy to solve the lack of training data. Then we introduce an instance guider to feed the color of the instance. To achieve accurate color matching, we present fine-grained color matching with edge loss to enhance visual quality. Equipped with the proposed modules, MagicColor enables automatically transforming sketches into vividly-colored images with accurate consistency and multi-instance control. Experiments on our collected datasets show that our model outperforms existing methods regarding chromatic precision. Specifically, our model critically automates the colorization process with zero manual adjustments, so novice users can produce stylistically consistent artwork by providing reference instances and the original line art. Our code and additional details are available at https://yinhan-zhang.github.io/color.
Figure 1: Visual results of MagicColor. Given a set of references and corresponding line art images, our approach enables coloring sketches while maintaining consistency across multiple instances. Compared to traditional methods, our approach significantly improves coloring efficiency.
$\dagger$ Equal contribution.* Corresponding author.
Figure 2: Illustration of the workflow of multi-instance sketch colorization production. Previous methods can only achieve multi-instance sketch colorization step by step, which is time-consuming and inaccurate. In contrast, our method can color sketch while maintaining consistency, making multi-instance sketch colorization easier.
1 Introduction
The fast development of the digital cartoon industry has demonstrated widespread potential applications for generative artificial intelligence. Colorizing cartoon sketches, a crucial aspect within the broader realm of the cartoon industry, serves as an indispensable task. It not only heightens the visual allure but also enriches the narrative experience by effectively vividly communicating emotions and actions. Using the automatic pipeline can streamline production workflows, accelerate content creation, reduce labor demands, and meet the growing demand for cartoons.
In the traditional process of cartoon sketch colorization, artists first analyze the line art to grasp the character and story context. They then choose suitable color palettes and manually apply colors layer by layer with attention to detail. However, traditional methods have two drawbacks. The first one is that they are time-consuming as manual coloring requires great effort to match the original design. For instance, in an animation series, multiple identical objects in the same image need to be colored one by one, which is very inefficient. The other problem is inaccuracy since manual work is error-prone and different artists may have varied color interpretations, causing inconsistent results in large-scale projects. Our method slots into this workflow, supporting auto-colorization while preserving design fidelity and instance-level color consistency.
Cartoon sketch colorization has garnered considerable research interest due to its potential in digital art and animation production. Currently, user-guided methods [7, 96, 23, 65, 59, 100, 5, 27, 3] rely on explicit human inputs such as color points, scribbles, or textual descriptions to direct the coloring process. Reference-based approaches [95, 102, 1, 20, 11] employ fused attention mechanisms to transfer colors from exemplar images while preserving structural coherence. While recent advances [58, 53] have improved automated colorization, three critical limitations persist: (1) Domain adaptation gap: existing colorization pipelines rely heavily on reference image fidelity, as pronounced structural mismatches between line art and exemplars frequently induce erroneous chromatic mappings. This dependency imposes an impractical requirement for near-isomorphic geometry between inputs, which rarely holds in unconstrained animation workflows where stylized sketches often diverge significantly from real-world references. (2) Instance-level control granularity: existing methods demonstrate a lack of fine-grained control over instance-specific attributes during the color transfer process. Discrepancies in character pose, proportion, or viewpoint between reference instances and target sketches often lead to distorted textures and the loss of important details. As a result, crucial features from the reference image may be significantly diminished during the colorization process. (3) Color consistency: achieving instance-aware color consistency is essential. The coloring of regions with non-closed lines or multiple objects often results in color bleeding, which compromises the accuracy and stability of instance control. This instability undermines the overall harmony of the image, which detracts from the viewer’s experience.
To overcome these challenges, we propose MagicColor, a novel framework that streamlines the colorization process. Our approach builds upon pre-trained diffusion model priors, capitalizing on their learned capacity to enforce visual consistency across generated outputs. The architectural cornerstones of our approach are articulated as follows. First, to tackle the misalignment between the reference instances and target line art, we incorporate an explicit reference mechanism that seamlessly infuses reference-derived color semantics and artistic styles into line art. In addition, we employ implicit latent control to ensure precise reference to each instance, significantly enhancing color accuracy and consistency. Second, to further improve the visual quality and color consistency of the output, we introduce edge loss and color matching. These methods compel the model to genuinely extract color information from the reference character design, thereby improving the accuracy of semantic correspondence and reducing the reliance on any color information that may inadvertently leak from the reference image. Third, our model adopts a two-stage, self-play training strategy, which addresses the challenges of limited multi-instance training data, and subsequently incorporates additional reference images to refine the colorization capability. By facilitating colorization across multiple instances, our model achieves impressive color consistency with minimal human intervention. Our approach achieves state-of-the-art results across both quantitative metrics and qualitative evaluations, outperforming prior art in animated content creation. We aim for a critical step toward fully automated, high-efficiency animation pipelines with guaranteed stylistic coherence. This method can also be extended beyond anime to the broader digital art and media fields. Our contributions can be summarized as follows:
  • We propose MagicColor, the first multi-instance coloring method to support multiple instances integration for sketch colorization in a single forward.
  • Technically, to solve the lack of multi-instance data, we design a two-stage, self-play training strategy. We also propose instance guider and pixel-level color matching with edge loss to enhance the color correspondence.
2 Related Work
2.1 Line Art Colorization
Line art colorization techniques strive to decode the link between semantics and color by leveraging large-scale datasets [98, 50, 51, 48, 49, 52, 94, 8, 64, 92, 93]. Researchers utilize a series of semantic modules such as classification [22], semantic segmentation [17, 103], and instance-aware information [70, 63] to enrich color vibrancy. These methods generally perform well when the object’s color has a strong semantic-based determinacy. However, when confronted with objects that exhibit a broad range of colors, they often produce lackluster results. In contrast, our proposed framework, equipped with an innovative imagination module, is engineered to transcend this constraint.
On the other hand, generative priors enshrined within pretrained Generative Adversarial Networks (GANs) [16, 73, 82] and Diffusion models have been cornerstones in pursuing photorealistic colorization. GANs are adept at both conditional and unconditional image synthesis. Specific conditioning factors, such as layout and semantic maps, are harnessed to fine-tune the image synthesis process [71, 72]. For example, the StyleGAN-family models have demonstrated remarkable prowess in generating high-resolution images without explicit conditioning [25, 26, 83]. Some models struggle to maintain the local spatial integrity of grayscale inputs but can use diverse color priors from pre-trained models  [63, 21, 13]. In our sketch colorization task, we aim to adapt these generative prior-based techniques. Our method can preserve sketch structures and semantics while using their color-generation abilities.
2.2 Visual Correspondence
In computer vision, visual correspondence aims to identify and match relevant features or points among different images. It is widely applied in tasks such as stereo vision and motion tracking. In the past, traditional methods [2, 46] relied on hand-designed features to establish corresponding relationships. Nowadays, deep learning methods [12, 31, 28, 9] obtain matching capabilities through supervised learning with labeled datasets. However, supervised learning faces significant scalability issues. Precise pixel-level annotations are not only time-consuming and labor-intensive but also costly. To address this, scholars have started to explore weakly-supervised or self-supervised visual correspondence models. For example, LightGlue [38] can match sparse local features across image pairs, through an adaptive mechanism. CoTracker [24] adopts a semi-supervised training method, by generating pseudo-labels using off-the-shelf models. DIFT [74] extracts features by diffusion models and can achieve pixel-level semantic point matching. Building upon the diffusion models’ prior correspondence knowledge, our framework enables reference-based colorization via semantic-aligned color mapping between line art and reference instances, without any structural modifications.
2.3 Reference-Based Image Colorization
A great deal of research has been committed to colorizing photographs using reference-based priors [18, 4, 91, 15, 75]. Initially, efforts focused on the transfer of chromatic information to the corresponding regions through luminance and texture alignment, with various low-level feature-based correspondence techniques developed for more precise local color transfer [32, 41, 7, 80, 53]. However, these methods are vulnerable to complex appearance variations of the same object as low-level features cannot capture semantic nuances [17, 99]. Line art colorization is notably different from natural image colorization [66, 84, 96, 89].
The Diffusion models have emerged as a powerful alternative. The Denoising Diffusion Probabilistic Model [47] and the Denoising Diffusion Implicit Model [68] paved the way for Latent Diffusion Models (LDMs) like Stable Diffusion [61], revolutionizing text-to-image generation. Building on LDMs, ControlNet  [97] uses task-specific conditions and multi-modal inputs [42, 86, 58] to control pre-trained diffusion models [85, 105, 14, 88, 104, 77, 87, 37]. Reference-based colorization [76, 54, 81, 56, 43, 44], guided by a user-provided reference image, has also become popular, with existing methods using strategies such as segmented graphs, active learning, and attention networks. AnimeDiffusion [6] and ColorizeDiffusion [90] introduce a diffusion-based reference-based framework for anime face colorization. Paint-by-Example [91] and ObjectStitch [69] leverage CLIP as their cross-modal encoder to extract instance-level visual-semantic embeddings, whereas AnyDoor [10] innovates by training on video sequences and adopting DINOv2 [55] for spatial-temporal feature learning. Despite these advancements, all frameworks focus on generic object categories, falling short of fine-grained part-level alignment required for intricate design tasks [78, 79].
3 Preliminaries
3.1 Latent Diffusion Model
As the core architecture of Stable Diffusion [61], Latent Diffusion Models (LDM) revolutionize text-to-image generation by executing diffusion-denoising processes in compressed latent space rather than raw pixel domain, enabling stable and efficient training. The pipeline begins with a Variational Autoencoder (VAE) projecting RGB images into low-dimensional latent codes, where semantic-guided diffusion sampling occurs under textual conditioning. Then, an UNet-based network incorporates self-attention and cross-attention mechanisms through UNet blocks to learn the reverse denoising process in latent space. Cross-attention establishes bidirectional interactions between text embeddings and visual features, ensuring prompt semantics are continuously infused. The whole training objective of the UNet can be written as:
$\mathcal{L}_{LDM}=\mathbb{E}_{t,\mathbf{z},\epsilon}\left[\left\|\epsilon-\epsilon_{\theta}\left(\sqrt{\alpha_{t}}\mathbf{z}+\sqrt{1-\alpha_{t}}\epsilon,c,t\right)\right\|^{2}\right],$ (1)
where $z$ notes the latent embedding of the training sample. $\epsilon_{\theta}$ and $\epsilon$ represent predicted noise by diffusion model and ground truth noise at corresponding timestep $t$, respectively. $c$ is the condition embedding involved in the generation and the co-efficient $\alpha_{t}$ remains consistent with that employed in vanilla diffusion models.
3.2 Reference Condition Injection
In the current scenario, reference instances frequently suffer from noise redundancy and semantic conflicts, which significantly impede the model’s ability to learn colors effectively. To address this issue, when presented with $N$ user-provided reference instances along with the line art input, our model employs a dual-branch condition injection strategy. This strategy aims to achieve semantic similarity and structural alignment with the input. First, we align multiple instances in a layer and input them into the reference net. Then, we encode with CLIP and apply reference attention mechanisms [18, 4, 45] to inject color and semantic info into the UNet. Also, the line art is injected into the UNet to enhance line-structure features. This lets the model learn global reference info evenly and avoid interference. Second, for instance-aware image generation, we utilize instance images $I_{i}$ embedding as latent control signals. Unlike other methods [63, 45], our paradigm supports the use of multiple instances, which enhances the model’s generalizability. Zero-shot customization is tough, so we use pre-trained vision models to extract target object identity. Previous studies used CLIP for target embedding [34, 60]. We use DINOv2 as the feature encoder [55] to get discriminative spatial identity. DINOv2, trained with patch-level objectives under random masking, has highly expressive features. Its output includes a $26\times 26\times 1024$ spatial embedding $s_{i}$ for patch-level features and a $1024$ dimensional global embedding $g_{i}$.
Figure 3: The overview of MagicColor pipeline. We combine a dual-UNet framework with an instance control module (ICM). During training, we use multiple instances to set the overall color accurately. The color matching enables the model to better align the colors of the target image with those of the reference instances precisely. The edge loss helps the model pay more attention to the high-frequency areas and edges, resulting in a more accurate and vivid colorization for each instance.
4 Method
Problem Definition. We first formulate the task of multi-instance sketch colorization with per-instance chromatic control, enabling precise mapping from multiple reference objects to corresponding line drawing instances via semantic correspondence. We can represent the problem in a more concise way using the following formula:
$I_{out}=(S,R,M),$ (2)
where $I_{out}$ is the output image to be generated. The condition $C$ is composed of a line art image $S$, a set of reference instances $R=\{[I_{1},\cdots,I_{i}]\mid i=1,\cdots,N\}$, and a set of instance masks $M=\{[M_{1},\cdots,M_{i}]\mid i=1,\cdots,N\}$. Each $I_{i}$ in $R$ corresponds to an instance, and each $M_{i}$ in $M$ is a binary mask indicating the reference spatial location of an instance.
Architecture Design. Owing to the high-fidelity detail demands in anime sketch colorization, the core challenge resides in designing encoders capable of spatially precise reference image parsing to extract sub-pixel-level visual cues. Inspired by recent studies [19, 35, 45] which demonstrated the effectiveness of leveraging an additional UNet architecture for this purpose, we introduce a reference net following a similar design. Additionally, the sketch and instance guider are designed based on ControlNet [97], which provides an efficient condition for injecting line art and instance-level information into the generative process. Furthermore, we employ DINOv2 [55] to encode images and train a Feed-Forward Network layer to extract features from the image embeddings. These extracted features are then integrated into the latent control signals.
4.1 Self-Play Training Strategy
We then introduce a two-stage, self-play training approach to solve the problem of lack of multi-instance training data and gradually boost the model’s performance to build an advanced model that can give high-quality coloring results.
Single-Reference Colorization Training. In the first stage, training starts by activating the reference net, UNet, and the sketch guider. In both two stages, we use color matching with edge loss for training. For each anime sequence, we perform random frame sampling without replacement: one frame is designated as the style reference exemplar, while the other serves as the raw input sketch for colorization. We use the whole reference image as conditional input and extract the line art from the original image. Given a reference and line art image pair, the model maps color semantics from the reference image to the line art by minimizing a carefully designed edge loss function, aiming to make the model learn basic color-semantic relationships, forming the basis for later multi-instance refinement.
Multi-Instance Refinement. In the second stage, we train the reference net, UNet, sketch guider, and instance guider to improve the model’s ability to handle multiple instances and make local coloring more accurate. To precisely control each instance in the references, we use the Segment Anything Model (SAM) [29, 101, 39] to extract instances separately and perform operations like random fusion, scaling, shuffling, and adding noise to instances in a layer as reference net input to enhance the model’s per-instance semantic perception for accurate colorization. Then, the model encodes all instances, aligning them in latent space, and starts coloring each instance from the sketch, considering their features and overall relationships. This process results in more accurate colorization, especially in areas with high color and semantic requirements.
4.2 Instance Control Module
Previous methods often resorted to directly imposing constraints within the reference attention mechanism for control purposes [18, 4, 19]. Nevertheless, when faced with the task of handling multiple instances as conditional input, this conventional approach proved to be extremely arduous. To uphold semantic correspondence and safeguard details, we align image embeddings in latent space, thereby empowering more precise and effective handling of the intricate input scenarios presented by multiple instances. Given target regions (instance masks), we align the bounding rectangle of the mask with the input DINO feature maps using box coordinates. Subsequently, we use the dense grids of the mask within this rectangle to interpolate features from the input DINO features. This process is similar to ROI alignment, but the resolution of ROI is flexible and follows the size of the mask. Finally, to further incorporate global information, we randomly drop 10% of the spatial embedding $s_{i}$ and replace it with the DINO global embedding $g_{i}$. As our target mask may have a different silhouette from the shape of the input image, injecting global embedding enhances the model’s generalizability and adaptability across diverse image conditions. The final ROI feature is then warped into the original region of the latent control signal:
$l_{c}=\sum_{i=1}^{n}Drop\_out(Interpolate(s_{i},M_{i},B_{i}),g_{i}),$ (3)
where $B_{i}$ is the bounding box of the target mask $M_{i}$. To establish an instance-aware colorization framework, we use a latent control signal, denoted as $l_{c}$, which is a latent feature with a size of $l_{c}\in\mathbb{R}^{C\times H^{\prime}\times W^{\prime}}$. As shown in Figure 4, our model with an instance control module achieves precise instance-level control to produce varied colorization outcomes. Although the training data is derived from anime datasets, the references can incorporate real-life images, highlighting the model’s adaptability across different domains and flexible utilization of diverse reference sources.
4.3 Structure-Content Supervision Enhancement
We present a comprehensive approach to enhance the performance of diffusion models in line art colorization. By introducing edge loss and color matching, we address crucial aspects of image quality, resulting in more visually appealing and perceptually accurate generated images.
Edge Loss. During the diffusion model’s training process, each pixel in an image contributes equally to the supervision process. The diffusion model’s original training objective is pixel-level mean squared error (MSE) loss. However, simultaneously, we aim to make the image content more consistent with human perception of image quality. To enhance the supervision over the high-frequency area and improve the generation quality, we propose the edge loss, which consists of the perceptual loss and the re-weighted edge loss. The formula for the perceptual loss is as follows:
$\mathcal{L}_{\text{perceptual}}=\frac{1}{N}\sum_{i=1}^{N}\left|\phi_{i}(z)-\phi_{i}(\hat{z})\right|_{2}^{2},$ (4)
where $\hat{z}$ is the prediction latent embedding obtained by decoding $\epsilon_{\theta}$, $N$ represents the number of feature layers used for calculation, and $\phi_{i}$ is the feature-extraction function corresponding to the $i$-th layer of the pre-trained neural network. $\lambda$ is a hyper-parameter that balances $\mathcal{L}_{LDM}$ and the perceptual loss $\mathcal{L}_{perceptual}$. Moreover, in scenes characterized by complex hierarchical structures and overlapping objects, the importance of object edges may differ significantly from that of the plain background pixels. We calculate the overall instances edge map, which effectively mitigates interference from the background as we focus solely on the edges of individual objects. Then, we apply structure foreground enhancement from SyntheOcc [33] to the edge pixels. Since we perform edge detection in the latent space with the size of the input images, we regard the edge map as a loss weight map $w$ to enhance structure supervision. Finally, our total loss can be written as:
$\mathcal{L}$ $=w\cdot\mathcal{L}_{LDM}+\lambda\cdot\mathcal{L}_{perceptual}$ (5)
Figure 4: Instance Control Ability. With the same line art and diverse reference instances, our method achieves precise instance-level control for varied colorization and all without needing extra guidance.
Figure 5: Inference Pipeline. When users input a line art alongside one or multiple reference instances, our model automatically fuses instances across different layers, enhances line trajectories in the line art, and then performs colorization.
Color Matching. We employ pre-trained diffusion models to identify corresponding points within real-world images. Our model, which has a UNet at its core, processes noisy images by cleaning them and extracting features crucial for establishing correspondences. Subsequently, we utilize a diffusion feature map to match pixels between two images. This is accomplished by determining the nearest neighbors and computing the similarity using cosine distance. To obtain pixel correspondences, we begin by extracting dense features from both images and matching them. Let $F$ be the feature map of an image $I$. For a pixel at position $p$, we obtain the feature vector $F(p)$ through bilinear interpolation. In terms of color matching, $C_{r}$ denotes the color feature of the reference image, and $C_{s}$ denotes the color feature of the source image. We obtain $C_{s}$ and $C_{r}$ by sampling features from the normalized source and target features. To quantify their similarity, we calculate the Euclidean distance ($D$) between $C_{r}$ and the flattened $C_{s}$. Given two feature vectors $v_{1}=C_{r}$ and $v_{2}=C_{s}$, the Euclidean distance is computed as follows:
$D=\sqrt{\sum_{k=1}^{f}(v_{1k}-v_{2k})^{2}},$ (6)
where $f$ represents the feature dimension. Subsequently, we find the nearest-neighbor indices. Finally, given a pixel $p_{1}$ in $v_{1}$, we find the corresponding pixel $p_{2}$ in $v_{2}$ as follows:
$p_{2}=\mathop{\arg\min}_{p}\ \ d\left(F_{1}\left(p_{1}\right),F_{2}(p)\right),$ (7)
where $d$ is the cosine distance. This approach transfers color-related information between the two images.
5 Experiments
5.1 Implementation Detail
Dataset. The dataset employed in this study is comprised of two key components: the anime video dataset and the image dataset. The data were preprocessed from Sakuga [57], ATD-12K [67] and manually collected from the internet. A total of 1,000 image pairs were randomly selected to form the test set and the remaining data constituted the training set. We then extracted their instances as references.
Baseline design. The backbone architecture for both the UNet and reference net is derived from Stable Diffusion 1.5 [62]. The sketch guider and instance guider are initialized with pre-trained ControlNet weights. The instance encoder is initialized using DINOv2.
Hyper-parameters. During training, we resized the height and weight of the input image to 512 and kept the original aspect ratio. Our model undergoes 100,000 optimization iterations with a batch size of 1. The learning rate is set to $1\times 10^{-5}$. The training phase takes around 7 days using 2 NVIDIA A800 80G GPUs. We perform a random horizon flip and random brightness adjustment for each reference image as data augmentation to simulate multi-view conditions.
5.2 Comparison
5.2.1 Qualitative Results
We compare our approach with previous sketch colorization methods. For GAN-based approaches, we consider RSIC [30] and SGA [36]. RSIC and SGA utilize GANs for colorization, each with a distinct architecture and training strategy. We observe that GAN-based methods tend to produce color incoordination. Moreover, when there are numerous color block regions in the image, color bleeding issues are likely to occur, where the color of one color block contaminates the adjacent color blocks. Regarding diffusion-based models, AnimeDiffusion [6], colorizeDiffusion [90], and MangaNinja[45] (without point guided) perform better in terms of color control. However, as shown in Figure 8, they still struggle to effectively transfer the colors from the reference instances to the line drawing sketch, especially for the colors of smaller details in the images. The root cause is that the models fail to fully learn the color correspondences between the original image and the reference instances.
Table 1: Quantitative comparisons with SOTA baselines.
FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
RSIC[36] 78.954 20.527 0.719 0.395
SGA[40] 101.384 18.771 0.679 0.436
AnimeDiffusion[6] 118.85 11.037 0.482 0.549
ColorizeDiffusion[90] 67.161 10.680 0.435 0.515
MangaNinja[45] 84.239 11.608 0.434 0.577
Ours 28.443 23.493 0.805 0.277
Figure 6: Ablations on each component. “w/o edge loss” indicates without the edge loss, “w/o color matching” indicates without the color matching, “w/o instance guider” indicates without the instance guider.
Figure 7: User study results. Quality: Visual fidelity of generated outputs, Similarity: Alignment with reference instances, Usability: Operational simplicity, Controllability: Consistency between references and result outputs.
5.2.2 Quantitative Results
To comprehensively evaluate the colorization ability of our model, we quantitatively compared our method with the state-of-the-art colorization method on our test set. The results are presented in Table LABEL:Quantitative_result. We did not consider background noise when evaluating the experimental results. Due to the limited resolution of most previous work, all measurements are performed at a resolution of $512\times 512$. We report four metrics of different methods. FID gauges visual similarity, with lower values indicating better quality. PSNR measures distortion, and higher values imply less distortion. SSIM assesses luminance, contrast, and structure similarity, with values closer to 1 showing better structural preservation. LPIPS measures perceptual similarity and lower values mean closer resemblance to the reference. By comparing our model with these GAN-based and diffusion-based methods using the established metrics, our approach demonstrates a significant advantage over previous methods.
Table 2: Quantitative results of ablation study.
FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
All 62.531 22.587 0.806 0.203
w/o Edge Loss 83.042 20.842 0.755 0.217
w/o Instance Guider 91.852 16.467 0.718 0.308
w/o Color Matching 88.573 20.832 0.749 0.226
Figure 8: Qualitative comparisons with existing methods. Given a line drawing and reference instances, our method demonstrates far more precise colorization and higher-quality results compared to other methods, effectively maintaining line-drawing structure and color consistency.
5.3 Ablation Study
Effectiveness of Edge Loss. To evaluate edge loss, we replace its edge loss with the original diffusion loss. Our method fails to preserve structural edge information, resulting in mismatched colors on the apple’s edges, as shown in the first row of Figure 6. When we eliminate the overall structure edge loss, the model struggles to maintain subtle edge features. Numerical evaluations in Table LABEL:ablation_study confirm the importance of structure edge loss for preserving image integrity.
Effectiveness of Color Matching. To assess our color-matching method, we remove the color matching component. This omission diminishes our model’s ability to retain color information and details from input references. For instance, in the second row of Figure 6, the colors of the bag and the character’s facial details are inconsistent.
Effectiveness of Instance Guider. To illustrate the effectiveness of our instance guider, we remove it during the experiment. Visual results in Figure 6 show that without this module, our model struggles to transfer instance-level color information from references effectively. As can be seen in the third row, the lack of an instance guider hampers the learning of instance-level information, leading to significant color loss and noise in the final results when relying solely on the reference net.
5.4 User Study
We developed a user interface to evaluate our proposed methods from user experiments. We invited 16 volunteers to experience our demonstration. Participants were required to perform reference-based colorization for all proposed models. The average testing time for each experimental set was more than 20 minutes. After the test was completed, we collected participants’ subjective evaluation scores in the following four dimensions: image quality, usability, reference similarity, and correlation between manipulated results. According to Figure 7, our method achieves the best performance in all four perspectives, especially in higher user satisfaction with image quality and control. Meanwhile, we can refine the proposed manipulation method further for better simplicity.
5.5 Limitations and Discussion
Our method has many advantages and potential applications but also faces several limitations. We summarize them as follows: (1) Flexible usage: When stylizing the same sketch with diverse reference images, our method retains the character’s identity. It adjusts details like lighting and background based on the reference styles. (2) Semantic awareness: For cartoon wallpapers or posters with multiple characters or objects, our approach uses multiple reference images to perform semantic-based colorization, ensuring each line art element gets colorized semantically. (3) Multi-object and occlusions: Single-subject line art can efficiently convert a sketch into a vivid, full-colored illustration, speeding up the production and enabling animators to test different color concepts quickly. However, in sketch images with many main objects or significant occlusions between them, the detailed inter-object colors may not be well maintained. (4) More explorations: The model can accurately color each element in multi-subject scenarios such as an anime battle scene or a forest scene while maintaining a harmonious overall color scheme.
6 Conclusion
This paper has presented MagicColor, a diffusion-based sketch colorization framework. Specifically, we introduce a multi-instance approach, a two-stage, self-play training strategy, and pixel-level color matching with edge loss. Our experiments demonstrated that MagicColor outperforms current methods in visual quality and style consistency, advancing the field of digital cartoon colorization. In future work, we plan to substantially augment the number of reference instances and enhance the model’s capacity to maintain semantic and color consistency. We also plan to release the source code for better collaboration between creative practitioners and AI researchers.

### v2

@@
-MagicColor: Multi-Instance Sketch Colorization
+Follow-Your-Color: Multi-Instance Sketch Colorization
 Abstract
-We present MagicColor, a diffusion-based framework for multi-instance sketch colorization.
+We present Follow-Your-Color, a diffusion-based framework for multi-instance sketch colorization.
 The production of multi-instance 2D line art colorization adheres to an industry-standard workflow, which consists of three crucial stages: the design of line art characters, the coloring of individual objects, and the refinement process.
@@
 Meanwhile, current generative methods fail to solve this task due to the challenge of multi-instance pair data collection.
-To tackle these challenges, we incorporate three technical designs to ensure precise character detail transcription and achieve multi-instance sketch colorization in a single forward.
-Specifically, we first propose the self-play training strategy to solve the lack of training data.
+To tackle these challenges, we incorporate three technical designs to ensure precise character detail transcription and achieve multi-instance sketch colorization in a single forward pass.
+Specifically, we first propose the self-play training strategy to address the lack of training data.
 Then we introduce an instance guider to feed the color of the instance.
 To achieve accurate color matching, we present fine-grained color matching with edge loss to enhance visual quality.
-Equipped with the proposed modules, MagicColor enables automatically transforming sketches into vividly-colored images with accurate consistency and multi-instance control.
+Equipped with the proposed modules, Follow-Your-Color enables automatically transforming sketches into vividly-colored images with accurate consistency and multi-instance control.
 Experiments on our collected datasets show that our model outperforms existing methods regarding chromatic precision.
@@
 Our code and additional details are available at https://yinhan-zhang.github.io/color.
-Figure 1: Visual results of MagicColor.
-Given a set of references and corresponding line art images, our approach enables coloring sketches while maintaining consistency across multiple instances.
+Figure 1: Results of Follow-Your-Color.
+Given a set of colored references, Follow-Your-Color can colorize a line art image while maintaining color consistency across multiple instances.
 Compared to traditional methods, our approach significantly improves coloring efficiency.
@@
 Previous methods can only achieve multi-instance sketch colorization step by step, which is time-consuming and inaccurate.
-In contrast, our method can color sketch while maintaining consistency, making multi-instance sketch colorization easier.
+In contrast, our method can color a sketch while maintaining consistency, making multi-instance sketch colorization easier.
 1 Introduction
@@
 Colorizing cartoon sketches, a crucial aspect within the broader realm of the cartoon industry, serves as an indispensable task.
-It not only heightens the visual allure but also enriches the narrative experience by effectively vividly communicating emotions and actions.
+It not only heightens the visual allure but also enriches the narrative experience by effectively and vividly communicating emotions and actions.
 Using the automatic pipeline can streamline production workflows, accelerate content creation, reduce labor demands, and meet the growing demand for cartoons.
@@
 However, traditional methods have two drawbacks.
-The first one is that they are time-consuming as manual coloring requires great effort to match the original design.
+The first one is that they are time-consuming, as manual coloring requires great effort to match the original design.
 For instance, in an animation series, multiple identical objects in the same image need to be colored one by one, which is very inefficient.
@@
 This instability undermines the overall harmony of the image, which detracts from the viewer’s experience.
-To overcome these challenges, we propose MagicColor, a novel framework that streamlines the colorization process.
+To overcome these challenges, we propose Follow-Your-Color, a novel framework that streamlines the colorization process.
 Our approach builds upon pre-trained diffusion model priors, capitalizing on their learned capacity to enforce visual consistency across generated outputs.
@@
 These methods compel the model to genuinely extract color information from the reference character design, thereby improving the accuracy of semantic correspondence and reducing the reliance on any color information that may inadvertently leak from the reference image.
-Third, our model adopts a two-stage, self-play training strategy, which addresses the challenges of limited multi-instance training data, and subsequently incorporates additional reference images to refine the colorization capability.
+Third, our model adopts a two-stage, self-play training strategy, which addresses the challenges of limited multi-instance training data and subsequently incorporates additional reference images to refine the colorization capability.
 By facilitating colorization across multiple instances, our model achieves impressive color consistency with minimal human intervention.
@@
 Our contributions can be summarized as follows:
-  • We propose MagicColor, the first multi-instance coloring method to support multiple instances integration for sketch colorization in a single forward.
+  • We propose Follow-Your-Color, the first multi-instance coloring method to support multiple instances integration for sketch colorization in a single forward pass.
   • Technically, to solve the lack of multi-instance data, we design a two-stage, self-play training strategy.
-We also propose instance guider and pixel-level color matching with edge loss to enhance the color correspondence.
+We also propose an instance guider and pixel-level color matching with edge loss to enhance the color correspondence.
 2 Related Work
@@
 To address this, scholars have started to explore weakly-supervised or self-supervised visual correspondence models.
-For example, LightGlue [38] can match sparse local features across image pairs, through an adaptive mechanism.
-CoTracker [24] adopts a semi-supervised training method, by generating pseudo-labels using off-the-shelf models.
+For example, LightGlue [38] can match sparse local features across image pairs through an adaptive mechanism.
+CoTracker [24] adopts a semi-supervised training method by generating pseudo-labels using off-the-shelf models.
 DIFT [73] extracts features by diffusion models and can achieve pixel-level semantic point matching.
@@
 Initially, efforts focused on the transfer of chromatic information to the corresponding regions through luminance and texture alignment, with various low-level feature-based correspondence techniques developed for more precise local color transfer [32, 40, 7, 79, 52].
-However, these methods are vulnerable to complex appearance variations of the same object as low-level features cannot capture semantic nuances [17, 99].
+However, these methods are vulnerable to complex appearance variations of the same object, as low-level features cannot capture semantic nuances [17, 98].
 Line art colorization is notably different from natural image colorization [65, 83, 95, 88].
@@
 3.1 Latent Diffusion Model
-As the core architecture of Stable Diffusion [61], Latent Diffusion Models (LDM) revolutionize text-to-image generation by executing diffusion-denoising processes in compressed latent space rather than raw pixel domain, enabling stable and efficient training.
+As the core architecture of Stable Diffusion [60], Latent Diffusion Models (LDM) revolutionize text-to-image generation by executing diffusion-denoising processes in a compressed latent space rather than the raw pixel domain, enabling stable and efficient training.
 The pipeline begins with a Variational Autoencoder (VAE) projecting RGB images into low-dimensional latent codes, where semantic-guided diffusion sampling occurs under textual conditioning.
-Then, an UNet-based network incorporates self-attention and cross-attention mechanisms through UNet blocks to learn the reverse denoising process in latent space.
+Then, a UNet-based network incorporates self-attention and cross-attention mechanisms through UNet blocks to learn the reverse denoising process in the latent space.
 Cross-attention establishes bidirectional interactions between text embeddings and visual features, ensuring prompt semantics are continuously infused.
@@
 where $z$ notes the latent embedding of the training sample.
-$\epsilon_{\theta}$ and $\epsilon$ represent predicted noise by diffusion model and ground truth noise at corresponding timestep $t$, respectively.
-$c$ is the condition embedding involved in the generation and the co-efficient $\alpha_{t}$ remains consistent with that employed in vanilla diffusion models.
+$\epsilon_{\theta}$ and $\epsilon$ represent predicted noise by the diffusion model and ground truth noise at corresponding timestep $t$, respectively.
+$c$ is the condition embedding involved in the generation, and the coefficient $\alpha_{t}$ remains consistent with that employed in vanilla diffusion models.
 3.2 Reference Condition Injection
-In the current scenario, reference instances frequently suffer from noise redundancy and semantic conflicts, which significantly impede the model’s ability to learn colors effectively.
+In the current scenario, reference instances frequently suffer from background noise, redundancy, and semantic conflicts (e.g., similar instances coexist in the same reference image, or the background contains complex textures), which significantly impede the model’s ability to learn colors effectively.
 To address this issue, when presented with $N$ user-provided reference instances along with the line art input, our model employs a dual-branch condition injection strategy.
@@
 Unlike other methods [62, 44], our paradigm supports the use of multiple instances, which enhances the model’s generalizability.
-Zero-shot customization is tough, so we use pre-trained vision models to extract target object identity.
+Zero-shot customization is tough, so we use pre-trained vision models to extract the target object identity.
 Previous studies used CLIP for target embedding [34, 59].
@@
 Its output includes a $26\times 26\times 1024$ spatial embedding $s_{i}$ for patch-level features and a $1024$ dimensional global embedding $g_{i}$.
-Figure 3: The overview of MagicColor pipeline.
+Figure 3: Overview of the Follow-Your-Color pipeline.
 We combine a dual-UNet framework with an instance control module (ICM).
@@
 where $I_{out}$ is the output image to be generated.
-The condition $C$ is composed of a line art image $S$, a set of reference instances $R=\{[I_{1},\cdots,I_{i}]\mid i=1,\cdots,N\}$, and a set of instance masks $M=\{[M_{1},\cdots,M_{i}]\mid i=1,\cdots,N\}$.
+The condition is composed of a line art image $S$, a set of reference instances $R=\{[I_{1},\cdots,I_{i}]\mid i=1,\cdots,N\}$, and a set of instance masks $M=\{[M_{1},\cdots,M_{i}]\mid i=1,\cdots,N\}$.
 Each $I_{i}$ in $R$ corresponds to an instance, and each $M_{i}$ in $M$ is a binary mask indicating the reference spatial location of an instance.
 Architecture Design.
-Owing to the high-fidelity detail demands in anime sketch colorization, the core challenge resides in designing encoders capable of spatially precise reference image parsing to extract sub-pixel-level visual cues.
+Owing to the high-fidelity detail demands in anime sketch colorization, the core challenge resides in designing encoders capable of spatially precise reference image parsing to extract pixel-level visual cues.
 Inspired by recent studies [19, 35, 44] which demonstrated the effectiveness of leveraging an additional UNet architecture for this purpose, we introduce a reference net following a similar design.
@@
 4.1 Self-Play Training Strategy
-We then introduce a two-stage, self-play training approach to solve the problem of lack of multi-instance training data and gradually boost the model’s performance to build an advanced model that can give high-quality coloring results.
+We then introduce a two-stage, self-play training approach to solve the problem of a lack of multi-instance training data and gradually boost the model’s performance to build an advanced model that can give high-quality coloring results.
 Single-Reference Colorization Training.
 In the first stage, training starts by activating the reference net, UNet, and the sketch guider.
-In both two stages, we use color matching with edge loss for training.
+In both stages, we use color matching with edge loss for training.
 For each anime sequence, we perform random frame sampling without replacement: one frame is designated as the style reference exemplar, while the other serves as the raw input sketch for colorization.
-We use the whole reference image as conditional input and extract the line art from the original image.
+We use the whole reference image as a conditional input and extract the line art from the original image.
 Given a reference and line art image pair, the model maps color semantics from the reference image to the line art by minimizing a carefully designed edge loss function, aiming to make the model learn basic color-semantic relationships, forming the basis for later multi-instance refinement.
@@
 Nevertheless, when faced with the task of handling multiple instances as conditional input, this conventional approach proved to be extremely arduous.
-To uphold semantic correspondence and safeguard details, we align image embeddings in latent space, thereby empowering more precise and effective handling of the intricate input scenarios presented by multiple instances.
+To uphold semantic correspondence and safeguard details, we align image embeddings in the latent space, thereby empowering more precise and effective handling of the intricate input scenarios presented by multiple instances.
 Given target regions (instance masks), we align the bounding rectangle of the mask with the input DINO feature maps using box coordinates.
 Subsequently, we use the dense grids of the mask within this rectangle to interpolate features from the input DINO features.
-This process is similar to ROI alignment, but the resolution of ROI is flexible and follows the size of the mask.
+This process is similar to ROI alignment, but the resolution of the ROI is flexible and follows the size of the mask.
 Finally, to further incorporate global information, we randomly drop 10% of the spatial embedding $s_{i}$ and replace it with the DINO global embedding $g_{i}$.
@@
 Although the training data is derived from anime datasets, the references can incorporate real-life images, highlighting the model’s adaptability across different domains and flexible utilization of diverse reference sources.
+Table 3: Quantitative Comparisons Across Datasets
+Domain Dataset FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+Animation Hand-drawn Animation Hand-drawn Animation Hand-drawn Animation Hand-drawn
+RSIC(256x) 128.971 300.636 10.933 8.872 0.548 0.416 0.508 0.566
+SGA(256x) 124.732 306.881 10.021 7.253 0.521 0.414 0.454 0.533
+AnimeDiffusion(256x) 154.784 305.021 11.262 10.207 0.416 0.426 0.557 0.549
+ColorizeDiffusion(512x) 82.201 298.673 17.882 9.817 0.512 0.349 0.425 0.513
+MangaNinja(512x) 43.165 295.223 14.289 10.249 0.543 0.361 0.362 0.508
+Ours(512x) 28.953 251.301 23.751 10.551 0.783 0.434 0.201 0.431
 4.3 Structure-Content Supervision Enhancement
@@
 With the same line art and diverse reference instances, our method achieves precise instance-level control for varied colorization and all without needing extra guidance.
-Figure 5: Inference Pipeline.
-When users input a line art alongside one or multiple reference instances, our model automatically fuses instances across different layers, enhances line trajectories in the line art, and then performs colorization.
 Color Matching.
@@
 Our model, which has a UNet at its core, processes noisy images by cleaning them and extracting features crucial for establishing correspondences.
-Subsequently, we utilize a diffusion feature map to match pixels between two images.
+Two steps utilize a diffusion feature map to match pixels between two images.
+(1) Semantic Matching.
 This is accomplished by determining the nearest neighbors and computing the similarity using cosine distance.
 To obtain pixel correspondences, we begin by extracting dense features from both images and matching them.
-Let $F$ be the feature map of an image $I$.
+Let $F_{i}$ be the feature map of an image $i$.
 For a pixel at position $p$, we obtain the feature vector $F(p)$ through bilinear interpolation.
-In terms of color matching, $C_{r}$ denotes the color feature of the reference image, and $C_{s}$ denotes the color feature of the source image.
-We obtain $C_{s}$ and $C_{r}$ by sampling features from the normalized source and target features.
-To quantify their similarity, we calculate the Euclidean distance ($D$) between $C_{r}$ and the flattened $C_{s}$.
-Given two feature vectors $v_{1}=C_{r}$ and $v_{2}=C_{s}$, the Euclidean distance is computed as follows:
+In terms of color matching, $C_{reference}$ denotes the color feature of the reference image, and $C_{source}$ denotes the color feature of the source image.
+We obtain $C_{source}$ and $C_{reference}$ by sampling features from the normalized source and target features.
+To quantify their similarity, we calculate the Euclidean distance ($D$) between $C_{reference}$ and the flattened $C_{source}$.
+Given two feature vectors $v_{1}=C_{reference}$ and $v_{2}=C_{source}$, the Euclidean distance is computed as follows:
 $D=\sqrt{\sum_{k=1}^{f}(v_{1k}-v_{2k})^{2}},$ (6)
@@
 Subsequently, we find the nearest-neighbor indices.
-Finally, given a pixel $p_{1}$ in $v_{1}$, we find the corresponding pixel $p_{2}$ in $v_{2}$ as follows:
+(2) Feature injection.
+Given a pixel $p_{1}$ in $v_{1}$, we find the corresponding pixel $p_{2}$ in $v_{2}$ as follows:
 $p_{2}=\mathop{\arg\min}_{p}\ \ d\left(F_{1}\left(p_{1}\right),F_{2}(p)\right),$ (7)
 where $d$ is the cosine distance. This approach transfers color-related information between the two images.
+Figure 5: Qualitative comparisons with existing methods.
+Given a line drawing and multiple reference instances, our method demonstrates far more precise colorization and higher-quality results compared to other methods, effectively maintaining line-drawing structure and reference instances’ color consistency.
 5 Experiments
@@
 Dataset.
-The dataset employed in this study is comprised of two key components: the anime video dataset and the image dataset.
-The data were preprocessed from Sakuga [57], ATD-12K [67] and manually collected from the internet.
-A total of 1,000 image pairs were randomly selected to form the test set and the remaining data constituted the training set.
+The dataset employed in this study comprises two key components: the anime video dataset and the image dataset.
+The data were preprocessed from Sakuga [56], ATD-12K [66], and manually collected from the internet.
+A total of 1,000 image pairs were selected from animations to form the ablation test set, and the remaining data constituted the training set.
 We then extracted their instances as references.
@@
 Hyper-parameters.
-During training, we resized the height and weight of the input image to 512 and kept the original aspect ratio.
+During training, we resized the height and width of the input image to 512 and kept the original aspect ratio.
 Our model undergoes 100,000 optimization iterations with a batch size of 1.
@@
 5.2.1 Qualitative Results
+To achieve the same settings, we use a complete reference image as the reference condition for the comparative baseline methods, and our method employs instances extracted from the same reference image.
 We compare our approach with previous sketch colorization methods.
@@
 The root cause is that the models fail to fully learn the color correspondences between the original image and the reference instances.
-Table 1: Quantitative comparisons with SOTA baselines.
-FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-RSIC[36] 78.954 20.527 0.719 0.395
-SGA[40] 101.384 18.771 0.679 0.436
-AnimeDiffusion[6] 118.85 11.037 0.482 0.549
-ColorizeDiffusion[90] 67.161 10.680 0.435 0.515
-MangaNinja[45] 84.239 11.608 0.434 0.577
-Ours 28.443 23.493 0.805 0.277
-Figure 6: Ablations on each component.
-“w/o edge loss” indicates without the edge loss, “w/o color matching” indicates without the color matching, “w/o instance guider” indicates without the instance guider.
-Figure 7: User study results.
-Quality: Visual fidelity of generated outputs, Similarity: Alignment with reference instances, Usability: Operational simplicity, Controllability: Consistency between references and result outputs.
 5.2.2 Quantitative Results
 To comprehensively evaluate the colorization ability of our model, we quantitatively compared our method with the state-of-the-art colorization method on our test set.
-The results are presented in Table LABEL:Quantitative_result.
+We constructed an Animation Dataset (1,570 pairs) from two anime films (Your Name and Spirited Away) and a Hand-Drawn Dataset (100 pairs) collected from real-user sketches as a test set.
+The results are presented in Table 3.
 We did not consider background noise when evaluating the experimental results.
@@
 SSIM assesses luminance, contrast, and structure similarity, with values closer to 1 showing better structural preservation.
-LPIPS measures perceptual similarity and lower values mean closer resemblance to the reference.
+LPIPS measures perceptual similarity, and lower values mean closer resemblance to the reference.
 By comparing our model with these GAN-based and diffusion-based methods using the established metrics, our approach demonstrates a significant advantage over previous methods.
@@
 w/o Color Matching 88.573 20.832 0.749 0.226
-Figure 8: Qualitative comparisons with existing methods.
-Given a line drawing and reference instances, our method demonstrates far more precise colorization and higher-quality results compared to other methods, effectively maintaining line-drawing structure and color consistency.
 5.3 Ablation Study
@@
 Effectiveness of Color Matching.
-To assess our color-matching method, we remove the color matching component.
-This omission diminishes our model’s ability to retain color information and details from input references.
+Removing the color-matching component diminishes our model’s ability to retain color information and details from input references.
 For instance, in the second row of Figure 6, the colors of the bag and the character’s facial details are inconsistent.
@@
 As can be seen in the third row, the lack of an instance guider hampers the learning of instance-level information, leading to significant color loss and noise in the final results when relying solely on the reference net.
-5.4 User Study
-We developed a user interface to evaluate our proposed methods from user experiments.
-We invited 16 volunteers to experience our demonstration.
-Participants were required to perform reference-based colorization for all proposed models.
-The average testing time for each experimental set was more than 20 minutes.
-After the test was completed, we collected participants’ subjective evaluation scores in the following four dimensions: image quality, usability, reference similarity, and correlation between manipulated results.
-According to Figure 7, our method achieves the best performance in all four perspectives, especially in higher user satisfaction with image quality and control.
-Meanwhile, we can refine the proposed manipulation method further for better simplicity.
+Figure 6: Ablations on each component.
+“w/o edge loss” indicates without the edge loss, “w/o color matching” indicates without the color matching, “w/o instance guider” indicates without the instance guider.
 5.4 Limitations and Discussion
-Our method has many advantages and potential applications but also faces several limitations.
+Our method has many advantages and potential applications, but it also faces several limitations.
 We summarize them as follows: (1) Flexible usage: When stylizing the same sketch with diverse reference images, our method retains the character’s identity.
@@
 6 Conclusion
-This paper has presented MagicColor, a diffusion-based sketch colorization framework.
+This paper has presented Follow-Your-Color, a diffusion-based sketch colorization framework.
 Specifically, we introduce a multi-instance approach, a two-stage, self-play training strategy, and pixel-level color matching with edge loss.
-Our experiments demonstrated that MagicColor outperforms current methods in visual quality and style consistency, advancing the field of digital cartoon colorization.
+Our experiments demonstrated that Follow-Your-Color outperforms current methods in visual quality and style consistency, advancing the field of digital cartoon colorization.
 In future work, we plan to substantially augment the number of reference instances and enhance the model’s capacity to maintain semantic and color consistency.
-We also plan to release the source code for better collaboration between creative practitioners and AI researchers.
+7 Acknowledagment
+This work was partially supported by the Guangdong Provincial Key Lab of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things #2023B1212010007.
+The Research Grant Council of the Hong Kong Special Administrative Region under grant number 16203122.